3) Cathys cars acceleration is given by the following equation a = (4 m/s) A) Sketch Cathys motion on the three graphs bel
Answer #1

3) a= 4m83) as Vaasat » V = 41m) (b)=) V = 2 (/2ť? - and displacement x =vlat (C)= x= 2(m/5325 tällt = 20/12) X = 2/3 (753) +

If there's any doubt feel free to comment,thank u